OpenStudy (anonymous):

Four times a number plus five more than three times the number is 47.

OpenStudy (anonymous):

What is your question?

OpenStudy (anonymous):

Four times a number plus five more than three times the number is 47.

OpenStudy (anonymous):

The equation is 4x + (3x+5) = 47 To solve, combine like terms and isolate the variable.

OpenStudy (anonymous):

i still dnt get it

OpenStudy (anonymous):

Okay let the no. Be x.. So four times of it would be 4x. Now 3 times he no.+5=47... So ur equation would be 4x+(3x+5)=47.. Now solve it

OpenStudy (anonymous):

You combine the x's so that you have 7x + 5 = 47 You need to get rid of the 5 so you subtract it from both sides. You then have 7x = 42 Solve for x

OpenStudy (anonymous):

i think i got it ?

OpenStudy (anonymous):

What answer did you get?

OpenStudy (anonymous):

5 - 9n = 47; -5.77

OpenStudy (anonymous):

i dont think its .. right tho .. math is my worst subject

OpenStudy (anonymous):

No the answer is six and it has not been rounded off.. Okay shall we do this step by step? @deesmith ??

OpenStudy (anonymous):

Look u previously got an equation 4x+(3x+5)=47 correct? Now what you need to do is remove the brackets to have a simple equation4x+3x+5=47. Can u solve it now @deesmith ??

OpenStudy (anonymous):

Alright, here we go. Your original equation is 4x + 3x + 5 = 47 You need to get all the x's together so you add them 4x + 3x = 7x Your equation now looks like this, 7x + 5 = 47 Since you are trying to find x, you subtract the 5 from both sides. 5 - 5 =0 and 47 - 5 = 42 You now have 7x = 42 We need to find out what x is. Since it is multiplied by 7 (7x), we do the opposite, and divide 7x/7 = x (remember that what you do to one side, you must also do to the other) and 42/7 = 6 You end up with x = 6

OpenStudy (anonymous):

:) thanks ... dang your smart !!!

OpenStudy (anonymous):

math is my worst subject